OpenStudy (anonymous):

Find two positive numbers with product 200 such that the sum of one number and twice the second number is as small as possible. (calculus approach please)

OpenStudy (bahrom7893):

Find 2 positive numbers with product 200: \[x * y=200\]The sum of one number and twice the second number is as small as possible \[f(x)=x+2y\]

OpenStudy (bahrom7893):

Minimize f(x), so find f'(x), see where it's 0 or undefined, and all that.

OpenStudy (bahrom7893):

\[f(x)=x+2y=x+2*\frac{200}{x}\]

OpenStudy (bahrom7893):

\[f(x) = x + \frac{400}{x}=\frac{x*x+400}{x}=\frac{x^2+400}{x}\]

OpenStudy (bahrom7893):

What is the derivative of that function?

OpenStudy (anonymous):

Is it \[\frac{ x^2-400 }{ x^2 }\]???

OpenStudy (bahrom7893):

Can you post the steps?

OpenStudy (anonymous):

\[f(x)= \frac{ x^2 }{ x }+\frac{ 400 }{ x } \[=x+400x^{-1}\] \[f'(x)=1-400x ^{-2}\] \[=1-\frac{ 400 }{ x^2 }\] \[=\frac{ x^2-400 }{ x^2 }\]

OpenStudy (bahrom7893):

Correct, now, when is f'(x) equal to 0 or undefined?

OpenStudy (anonymous):

x=20?

OpenStudy (bahrom7893):

Actually, when x=+/- 20, and when x = 0. But since you know that the numbers are positive, and it can't be 0, you're right, your answer is 20.

OpenStudy (anonymous):

Oh okay, Thanks!

OpenStudy (bahrom7893):

np